# Key exchange protocol based on circulant matrix action over congruence-simple semiring

### v1

@@ Abstract
We present a new key exchange protocol based on circulant matrices acting on matrices over a congruence-simple semiring. We describe how to compute matrices with the necessary properties for the implementation of the protocol. Additionally, we provide an analysis of its computational cost and its security against known attacks.
@@ 1 Introduction
The foundational paper by [1] is considered the beginning of modern cryptography. In its original formulation, a cyclic group $\mathbb{Z}_{n}$ was used as the algebraic setting for the key exchange protocol. Later, a similar protocol based on elliptic curves was independently proposed by [3] and [5].
These protocols have become the standard in cryptography, and their use is widely spread. However, in [13], Shor presented a quantum algorithm capable of solving the discrete logarithm problem in both algebraic settings. Since the security of these cryptographic protocols relies on the hardness of this problem, a sufficiently large quantum computer would be able to break most of today’s key exchange protocols. As a result, the scientific community has initiated the search for new post-quantum cryptographic protocols, that is, protocols secure against quantum attacks.
@@ 2 Mathematical setting
Let $(R,+)$ a finite reducible semigroup with cardinal $n$, with $R=A\oplus B$. Entonces $1<\lvert A\rvert,\lvert B\rvert<n$.

### v2

@@ Abstract
We present a new key-exchange protocol based on circulant matrices acting on matrices over a congruence-simple semiring. We describe how to compute matrices with the necessary properties for the implementation of the protocol. In addition, we provide an analysis of its computational cost and its security against known attacks.
@@ 1 Introduction
The foundational paper by [1] is considered to be the beginning of modern cryptography. In its original formulation, a cyclic group $\mathbb{Z}_{n}$ was used as the algebraic setting for the key exchange protocol. Later, a similar protocol based on elliptic curves was independently proposed by [3] and [5].
These protocols have become the standard in cryptography, and their use is widely spread. However, in [13], Shor presented a quantum algorithm capable of solving the discrete logarithm problem in both algebraic settings. Since the security of these cryptographic protocols is based on the hardness of this problem, a sufficiently large quantum computer would be able to break most of today’s key exchange protocols. As a result, the scientific community has initiated the search for new post-quantum cryptographic protocols, that is, protocols secure against quantum attacks.
@@ 2 Mathematical setting
Let $(R,+)$ a finite reducible semigroup with cardinal $n$, with $R=A\oplus B$. Then $1<\lvert A\rvert,\lvert B\rvert<n$.